# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

### v2

@@ 4.3.2. Prior-Posterior Learning
Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned. In order to retain the codebook’s categorical properties for factors, we add the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020) to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively. Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
Due to stock market’s inherent volatility and complexit, it is difficult to bridge the gap between the noisy market and effective factor model (Duan et al., 2022). Therefore, we utilize the prior-posterior structure to predict future returns and optimize latent factor embeddings. In the training stage, the posterior layer predict posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
@@
The return predictor uses factor expected return $\mathbf{z}$ to calculate future returns:
@@ 5. Experiment
5.1. Experiment Settings
    where $T_{test}$ is the number of trading days of the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ are the true and predicted ranks of stocks on the trading day $s$.
Baselines. We compare STORM with nine methods across one prediction task and two downstream tasks. The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models. In the prediction task and PM task, we compare STORM with ML & DL-based and factor models. In the AT task, we compare STORM with ML & DL-based and RL-based methods. We use the labels P, PM, and AT to indicate the applicable task scope for the following baselines. The following will provide a brief introduction to each method:
  • Market
    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable. (P, PM, AT)
  • ML&DL-based
    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals. (P, PM)
    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions. (P, PM)
    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts. (P, PM)
  • RL-based
    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces. (PM, AT)
    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency. (PM, AT)
    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data. (PM, AT)
  • Factor-model
    • CAFactor is a state-of-the-art model that learns latent factors depending on asset characteristics. (P, PM)
    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors. (P, PM)
    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise. (P, PM)
@@ 5.1. Experiment Settings
STORM 0.137 1.472 0.501 0.134 1.459 0.363 0.239 1.903 1.263 0.099 1.339 0.433 0.250 1.954 0.724
± 0.015 ± 0.058 ± 0.035 ± 0.039 ± 0.151 ± 0.072 ± 0.033 ± 0.150 ± 0.224 ± 0.085 ± 0.321 ± 0.289 ± 0.034 ± 0.163 ± 0.086
STORM-w/o-TS 0.124 1.428 0.714 0.127 1.437 0.627 0.126 1.428 0.703 -0.001 1.000 0.099 0.103 1.341 0.360
± 0.062 ± 0.229 ± 0.003 ± 0.058 ± 0.214 ± 0.187 0± 0.022 ± 0.085 ± 0.091 ± 0.049 ± 0.152 ± 0.171 ± 0.017 ± 0.062 ± 0.065
STORM-w/o-CS 0.059 1.187 0.309 0.103 1.347 0.350 0.149 1.525 0.908 -0.031 0.913 0.048 0.148 1.513 0.445
± 0.021 ± 0.070 ± 0.061 ± 0.043 ± 0.154 ± 0.106 ± 0.095 ± 0.400 ± 0.450 ± 0.053 ± 0.156 ± 0.109 ± 0.03 ± 0.117 ± 0.040
Improvement(%) -6.803 -2.452 35.227 0.752 5.419 2.117 5.286 15.054 7.949 25.316 3.798 -5.459 -6.367 -5.467 -22.067
@@
STORM 0.913 0.309 0.017 0.983 0.389 0.015 1.843 0.162 0.011 0.700 0.326 0.019 1.441 0.312 0.016
± 0.068 ± $0$ $0$ 0.206 ± 0.008 ± $0$ 0.386 ± 0.026 ± 0.002 ± 0.336 ± 0.080 ± 0.002 ± 0.142 ± 0.058 ± 0.002
STORM-w/o-TS 0.943 0.250 0.014 1.587 0.194 0.011 1.223 0.174 0.010 0.203 0.418 0.018 0.800 0.336 0.015
± 0.354 ± 0.062 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.236 ± 0.014 ± 0.001 ± 0.284 ± 0.095 ± 0.002 ± 0.063 ± 0.044 ± 0.002
STORM-w/o-CS 0.605 0.246 0.013 0.858 0.327 0.013 1.426 0.157 0.010 0.112 0.491 0.020 0.959 0.363 0.017
± 0.114 ± 0.115 ± 0.004 ± 0.236 ± 0.064 ± 0.001 ± 0.509 ± 0.046 ± 0.004 ± 0.231 ± 0.161 ± 0.005 ± 0.092 ± 0.033 ± 0.001
Improvement(%) -10.531 -2.929 -30.000 19.144 8.491 -10.000 1.153 $0$11.111 14.943 -118.792 -80.000 -13.816 -71.429 -36.364
@@
5.2. Implement Details
We implement all experiments by using PyTorch and conduct the experiments on two 80GB memory NVIDIA RTX H100 GPUs. The code is provided in the repository  [^4] https://github.com/DVampire/Storm. The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ denotes the batch size. $N$ represents the number of stocks, 28 for the DJ30 index and 408 for the SP500 index. The dimension $W=64$ corresponds to the number of historical data days, and $D=152$ represents the total number of features, including the OHLC data, technical indicators, and temporal information.
For the time-series module, we use an 8-day period for each stock as a patch, meaning the patch size is $(8,1,D)$. In contrast, for the cross-sectional module, we use $N$ stocks for each day as a patch, so the patch size is $(1,N,D)$. For both the encoders and decoders, we employ stacked transformer blocks. Each encoder consists of 4 layers of blocks with 4 attention heads, while each decoder consists of 2 layers of blocks with 8 attention heads. The size of both codebooks is selected from the set {256, 512, 1024}, with 512 yielding the best results. A consistent embedding dimension of 256 is employed across the encoders, decoders, and codebooks.
5.3. Performance on Factor Quality (for RQ1)
To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes. In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets. Specifically, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, indicating the learned latent factors are significantly effective in capturing stock price trends.
The diversity of factors is crucial in financial investment, as it ensures the model captures the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor. In the codebook, each embedding represents a cluster center for a category of factor embeddings. To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data. For clarity, we group every five indices into one interval, as shown in Figure 3(a). The results indicate that the distribution of factor usage is relatively even, demonstrating that the model has achieved good diversity in the factor selection process.
@@ 5.3. Performance on Factor Quality (for RQ1)
We experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality. If the codebook is too large, it may lead to sparse factor representation; if too small, it may result in insufficient factor differentiation and information redundancy. As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance.
5.4. Performance on downstream tasks (for RQ2)
We use three types of metrics: profit, risk-adjusted profit, and risk to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading. For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 10.45% over the best baseline methods on three stocks in the AT task. For risk resistance, despite the test periods for both datasets, including the COVID-19 event, STORM still achieves the best AVO value, surpassing all comparison methods in the PM task. Moreover, considering the return and risk simultaneously, STORM performs the best on ASR with an average improvement of 58.64% over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task. Such an outstanding ability to balance return and risk is significant in real investment practice.
5.5. Notations
We provide the main notations used in the paper.
@@ 5.5. Notations
Table 4. Notations in the paper.
5.6. Ablation Study (for RQ3)
6. Conclusions and Future Work

### v3

@@ 4.3.2. Prior-Posterior Learning
Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned. In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively. Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (Duan et al., 2022). Therefore, we introduce the prior-posterior structure to predict future returns and optimize latent factor embeddings. In the training stage, the posterior layer estimates posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
@@
The return predictor computes future returns based on the factor expected return $\mathbf{z}$:
@@
The predicted returns $\hat{\mathbf{y}}$ can be directly integrated into portfolio optimization, algorithmic trading strategies, enabling dynamic rebalancing and improved risk management.
@@ 5. Experiment
5.1. Experiment Settings
    where $T_{test}$ denotes the number of trading days in the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ represent the true and predicted ranks of stocks on the trading day $s$, respectively.
Baselines. As shown in Table 2, we compare STORM with nine methods across one prediction task and two downstream tasks. The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models. We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task. Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks. The following will provide a brief introduction to each method:
  • Market
    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable.
  • ML&DL-based
    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions.
    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts.
  • RL-based
    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces.
    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency.
    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data.
  • Factor-model
    • CAFactor (Gu et al., 2021) is a state-of-the-art model that learns latent factors depending on asset characteristics.
    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise.
@@ 5.1. Experiment Settings
Task Market ML & DL-based RL-based Factor model Ours
B&H LGBM (Yang et al., 2020) LSTM (Yu and Li, 2018) Transformer (Yang et al., 2020) SAC (Haarnoja et al., 2018) PPO (Schulman et al., 2017) DQN (Mnih et al., 2013) CAFactor (Gu et al., 2021) FactorVAE (Duan et al., 2022) HireVAE (Wei et al., 2023) STORM
Prediction
Portfolio Management
Algorithmic Trading
  • * indicates that the model is applicable, while signifies that it is NOT applicable to the corresponding task.
Table 2. Comparison of baselines across different tasks
Implementation Details. All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each. The implementations of the ML & DL methods are based on Qlib (Yang et al., 2020). The input data is organized into batches with dimensions corresponding to historical days, stocks, and features, ensuring compatibility across tasks. As for other baselines, we adopt the default settings in their public implementations. Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported. Detailed information regarding the network architecture of STORM, as well as the optimizers and hyperparameters used for the two downstream tasks, is provided in Appendix B.
@@
STORM 0.229 1.857 0.750 0.174 1.621 0.559 0.236 1.893 1.470 0.173 1.625 0.773 0.290 2.154 1.216
± 0.033 ± 0.154 ± 0.066 ± 0.032 ± 0.133 ± 0.081 ± 0.039 ± 0.184 ± 0.445 ± 0.067 ± 0.284 ± 0.293 ± 0.052 ± 0.262 ± 0.597
STORM-w/o-TS 0.199 1.730 0.766 0.127 1.437 0.627 0.154 1.536 1.294 0.106 1.212 0.267 0.254 1.979 0.964
± 0.059 ± 0.260 ± 0.274 ± 0.058 ± 0.214 ± 0.187 ± 0.023 ± 0.089 ± 1.012 ± 0.120 ± 0.145 ± 0.135 ± 0.060 ± 0.093 ± 0.503
STORM-w/o-CS 0.148 1.521 0.641 0.103 1.347 0.350 0.152 1.534 0.814 0.136 1.474 0.541 0.181 1.650 0.648
± 0.069 ± 0.283 ± 0.184 ± 0.043 ± 0.154 ± 0.106 ± 0.054 ± 0.219 ± 0.352 ± 0.082 ± 0.300 ± 0.328 ± 0.051 ± 0.221 ± 0.361
Improvement(%) 55.782 23.062 45.076 30.827 17.124 2.117 3.965 14.4501 20.408 118.987 26.969 66.594 8.614 4.159 30.893
@@
STORM 1.346 0.282 0.016 1.245 0.299 0.014 2.107 0.137 0.009 1.189 0.227 0.015 1.672 0.210 0.015
± 0.162 ± $0$ 0.001 ± 0.106 ± 0.016 ± 0.001 ± 0.302 ± 0.033 ± 0.001 ± 0.386 ± 0.055 ± 0.002 ± 0.201 ± 0.067 ± 0.002
STORM-w/o-TS 1.199 0.262 0.017 1.587 0.194 0.011 1.868 0.131 0.008 0.580 0.344 0.017 1.553 0.244 0.015
± 0.300 ± 0.052 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.956 ± 0.077 ± 0.003 ± 0.229 ± 0.061 ± $0$ 0.187 ± 0.093 ± 0.004
STORM-w/o-CS 1.027 0.229 0.015 0.858 0.327 0.013 1.397 0.177 0.010 0.914 0.312 0.018 1.401 0.295 0.016
± 0.326 ± 0.030 ± 0.001 ± 0.236 ± 0.064 ± 0.001 ± 0.405 ± 0.028 ± 0.001 ± 0.397 ± 0.136 ± 0.004 ± 0.224 ± 0.090 ± 0.002
Improvement(%) 27.704 4.184 - 19.144 8.491 - 15.642 16.561 11.111 95.238 - - 2.076 - -
@@
5.2. Factor Quality (for RQ1)
To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes. In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets. Specifically, in terms of RankIC and RankICIR, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, as shown in Table  1. These results highlight the learned latent factors are significantly effective in capturing stock price trends.
The diversity of factors is essential in financial investment, as it allows the model to capture the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor. In the codebook, each embedding represents a cluster center for a category of factor embeddings. To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data. For clarity, we group every five indices into one interval, as illustrated in Figure 3(a). The results indicates a relatively even distribution of factor usage, highlighting the model’s effectiveness to maintain diversity during the factor selection process.
Furthermore, we experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality. A larger codebook may lead to sparse factor representations, reducing the effectiveness of factor utilization, while a smaller codebook could result in insufficient factor differentiation and information redundancy. Following standard practices, we evaluate five different codebook sizes, including 128, 256, 512, 1024, and 2048. As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance, striking a balance between expressiveness and efficiency.
@@ 5.2. Factor Quality (for RQ1)
5.3. Downstream Task Performance (for RQ2)
We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading.
For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task. These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance. For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event. Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound. Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 58.64% and 166.83% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task. Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
5.4. Ablation Study (for RQ3)
5.5. Discussion on Performance
Efficiency. STORM achieves remarkable efficiency in real-world trading scenarios. Unlike high-frequency trading systems requiring ultra-low latency and specialized hardware, our daily trading strategy operates efficiently on standard CPU hardware. This is enabled by a compact model architecture characterized by: i) a minimalist design with only a few transformer blocks (see Appendix B), and ii) a lightweight parameterization comprising 15.2 million FP32 parameters on DJ30. This design ensures seamless deployment on edge devices while supporting concurrent market analysis and trading.
Effectiveness. STORM exhibits exceptional performance compared to baseline methods, showcasing versatility across tasks including PM and AT. Unlike baselines that often excel in specific tasks but lack generalizability, STORM consistently delivers robust overall results. Moreover, STORM can balance returns and risks by emphasizing comprehensive investment performance over individual financial metrics. Collectively, these attributes substantiate that the additional computational cost incurred by STORM is well justified by its robustness and adaptability across diverse tasks and datasets.
Generalization. STORM demonstrates robust generalizability by achieving strong performance across markets with varying scales and distinct characteristics. It excels on both the DJ30, which reflects large-cap industrial blue chips, and the SP500, representing a broader market across diverse industries. Furthermore, its flexible encoder-decoder architecture allows the integration of additional layers such as Mamba  (Gu and Dao, 2023; Dao and Gu, 2024) and ensures applicability to diverse trading environments and asset classes.
6. Conclusions and Future Work
Appendix A Notations
We provide the main notations in Table A.5 used in the paper.
@@ Appendix A Notations
Table A.5. Notations in the paper.
Appendix B Implement Details
We implement all experiments by using PyTorch and conduct the experiments on two 80GB memory NVIDIA RTX H100 GPUs. The code is provided in the repository  [^4] https://anonymous.4open.science/r/Storm-2823. The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ denotes the batch size. $N$ represents the number of stocks, 28 for the DJ30 index and 408 for the SP500 index. The dimension $W=64$ corresponds to the number of historical data days, and $D=152$ represents the total number of features, including the OHLC data, technical indicators, and temporal information.
For the time-series module, we use an 4-day period for each stock as a patch, meaning the patch size is $(4,1,D)$. In contrast, for the cross-sectional module, we use $N$ stocks for each day as a patch, so the patch size is $(1,N,D)$. For both the encoders and decoders, we employ stacked transformer blocks. Each encoder consists of 4 layers of blocks with 4 attention heads, while each decoder consists of 2 layers of blocks with 8 attention heads. The size of both codebooks is selected from the set {128, 256, 512, 1024, 2048}, with 512 yielding the best results. A consistent embedding dimension of 256 is employed across the encoders, decoders, and codebooks.